OpenStudy (anonymous):

Solve. 8/5=1200/5(t−5)^2

sam (.sam.):

\[\frac{8}{5}=\frac{1200}{5 (t-5)^2} ?\]

OpenStudy (anonymous):

10+2\[10+2\sqrt{6} or 10 - 2\sqrt{6}\]

OpenStudy (anonymous):

Bring the denominator over. so you have  now you're just left with 8(t-5)^2 = 1200 because the the two 5's cancel each other out. From here divide both sides by 8 so you now have (t - 5)^2 = (1200/8) (t-5)^2 = 150 take the square root of both sides t-5 = sqrt(150) add 5 to both sides. t = sqrt(150) + 5 which is roughly 17.24